OpenStudy (anonymous):

The holy trinity girls’ basketball team is to be chosen randomly from fifteen Grade 12 students and ten Grade 11 students trying out for the team. The team will consist of 9 players. Determine the probability, to the nearest thousandth, that exactly three are grade 11 students

OpenStudy (anonymous):

can u draw out the work so i cn understnad it better

OpenStudy (anonymous):

G11 = 10, G12 = 15, Team = 25, (10/25) = 2/5 (2/5)(2/5)(2/5) = 8/125 probability of 3 g11 girls

OpenStudy (anonymous):

Grade 11 Students = 10 Grade 12 Students = 15 25 Students total Basketball Team = 9 players (10/25) chance of ONE Grade 11 student making the team thus (10/25)(10/25)(10/25) chance of three grade 11 students making the team

OpenStudy (anonymous):

okay thankyou. can u help me witha nothr question

OpenStudy (mathmate):

I suggest you solve this using combinatorics. Number of ways to make a team of 3 from 10 grade 11 students = C(10,3) Number of ways to make a team of 6 from 15 grade 12 students = C(15,6) Total number of ways to make a team of 9 from 25 students = C(25,9) Therefore probability that there are exactly 3 from grade 11 is: C(10,3)*C(15,6)/C(25,9) = 2184/7429 (about 29.4%).